A bus slows down with a constant acceleration. It's speed goes from 24 m/s to 16 m/s as it travels 50 m.

a) What far does it continue to roll before stopping?

b) How long does it take for it to stop from the moment where the velocity is 24 m/s?


1
Expert's answer
2021-10-18T11:13:07-0400

a)


"24^2-16^2=2(50)a\\a=3.2\frac{m}{s^2}\\16^2=2(3.2)s\\s=40\ m"

b)


"t=\frac{24}{3.2}=7.5\ s"